# VisionLLM v2: An End-to-End Generalist Multimodal Large Language Model for Hundreds of Vision-Language Tasks

### v2

@@ 12. Can you assist me in generating masks that isolate each category belonging to <class> in this image?
1. Can you examine the image and pinpoint the keypoint locations of the <class>? 2. Could you analyze the picture and determine the keypoint placement of the <class>? 3. Please inspect the image and locate the keypoints for <class>. 4. Can you evaluate the photo and identify where the keypoints of <class> are situated? 5. Look at the image and detect the keypoint positions of the <class>. 6. Please analyze this image and find the keypoints of <class>. 7. Can you check the image and show me where the keypoints of <class> are located? 8. Please find the exact keypoint position of the <class>. 9. Please observe the photo and identify the keypoint locations of the <class>. 10. Can you review the image and point out the keypoints of <class>?
@@
1. Give me a concise description of the image. Answer the question and localize each object. 2. Please briefly summarize the content of this image. Answer the question and localize each object. 3. What does this picture show? Please summarize briefly. Answer the question and localize each object. 4. Can you give me a quick overview of what’s depicted in this image? Answer the question and localize each object. 5. Could you describe the key elements in this photograph? Answer the question and localize each object. 6. Offer a brief explanation of what this image represents. Answer the question and localize each object. 7. Sum up the contents of this picture in one or two sentences. Answer the question and localize each object. 8. What is the main content in this image? Answer the question and localize each object. 9. Provide a brief caption for the picture. Answer the question and localize each object. 10. Could you give me a short description of the image? Answer the question and localize each object.
@@
1. Can you examine the image and segment the corresponding objects denoted as <regions>?
2. Where are the objects marked by <regions> in the image? Could you help me segment these objects?
3. Could you please segment all the corresponding objects according to the visual prompt as <regions>?
4. Can you help me draw the instance segmentation masks of <regions> in the picture?
5. Please help me find all the objects shown as <regions> and segment them.
6. I’d like to know the objects outlined by <regions>. Please help me draw their masks.
7. Given the <regions>, I need your help to segment the corresponding object masks.
8. Examine the image and identify all the objects that belong to the provided <regions>.
9. I’m interested in the objects labeled as <regions>. Could you please draw their instance masks?
10. There are some regions represented by <regions>. I need your assistance to find their corresponding objects.
@@ 4. Can you help me draw the instance segmentation masks of <regions> in the picture?
1. Generate image with caption: <caption>.
2. Can you give me the image with caption: <caption>.
3. Help me to generate this image: <caption>.
4. Generate the image according to the caption: <caption>.
5. According to the caption, generate the image: <caption>.
6. An image with caption: <caption>.
7. Can you visualize this caption: <caption>.
8. Create an image based on this caption: <caption>.
9. Generate a visual representation for this caption: <caption>.
10. Provide me with an image corresponding to this caption: <caption>.
11. Craft an image with the following caption: <caption>.
12. Generate an image accompanied by this caption: <caption>.
13. Turn this caption into an image: <caption>.
14. Generate an image reflecting this caption: <caption>.
15. Translate this caption into a visual representation: <caption>.
16. Produce an image that matches this caption: <caption>.
17. Create an image in line with this caption: <caption>.
18. Generate an image to illustrate this caption: <caption>.
19. Construct an image based on the given caption: <caption>.
20. Give me an image associated with this caption: <caption>.

### v3

@@ 12. Can you assist me in generating masks that isolate each category belonging to <class> in this image?
1. Can you examine the image and pinpoint the keypoint locations of the <class>?
2. Could you analyze the picture and determine the keypoint placement of the <class>?
3. Please inspect the image and locate the keypoints for <class>.
4. Can you evaluate the photo and identify where the keypoints of <class> are situated?
5. Look at the image and detect the keypoint positions of the <class>.
6. Please analyze this image and find the keypoints of <class>.
7. Can you check the image and show me where the keypoints of <class> are located?
8. Please find the exact keypoint position of the <class>.
9. Please observe the photo and identify the keypoint locations of the <class>.
10. Can you review the image and point out the keypoints of <class>?
@@
1. Give me a concise description of the image. Answer the question and localize each object.
2. Please briefly summarize the content of this image. Answer the question and localize each object.
3. What does this picture show? Please summarize briefly. Answer the question and localize each object.
4. Can you give me a quick overview of what’s depicted in this image? Answer the question and localize each object.
5. Could you describe the key elements in this photograph? Answer the question and localize each object.
6. Offer a brief explanation of what this image represents. Answer the question and localize each object.
7. Sum up the contents of this picture in one or two sentences. Answer the question and localize each object.
8. What is the main content in this image? Answer the question and localize each object.
9. Provide a brief caption for the picture. Answer the question and localize each object.
10. Could you give me a short description of the image? Answer the question and localize each object.
@@
1. Can you examine the image and segment the corresponding objects denoted as <regions>? 2. Where are the objects marked by <regions> in the image? Could you help me segment these objects? 3. Could you please segment all the corresponding objects according to the visual prompt as <regions>? 4. Can you help me draw the instance segmentation masks of <regions> in the picture? 5. Please help me find all the objects shown as <regions> and segment them. 6. I’d like to know the objects outlined by <regions>. Please help me draw their masks. 7. Given the <regions>, I need your help to segment the corresponding object masks. 8. Examine the image and identify all the objects that belong to the provided <regions>. 9. I’m interested in the objects labeled as <regions>. Could you please draw their instance masks? 10. There are some regions represented by <regions>. I need your assistance to find their corresponding objects.
@@
1. Generate image with caption: <caption>. 2. Can you give me the image with caption: <caption>. 3. Help me to generate this image: <caption>. 4. Generate the image according to the caption: <caption>. 5. According to the caption, generate the image: <caption>. 6. An image with caption: <caption>. 7. Can you visualize this caption: <caption>. 8. Create an image based on this caption: <caption>. 9. Generate a visual representation for this caption: <caption>. 10. Provide me with an image corresponding to this caption: <caption>. 11. Craft an image with the following caption: <caption>. 12. Generate an image accompanied by this caption: <caption>. 13. Turn this caption into an image: <caption>. 14. Generate an image reflecting this caption: <caption>. 15. Translate this caption into a visual representation: <caption>. 16. Produce an image that matches this caption: <caption>. 17. Create an image in line with this caption: <caption>. 18. Generate an image to illustrate this caption: <caption>. 19. Construct an image based on the given caption: <caption>. 20. Give me an image associated with this caption: <caption>.